# Data-Free Continual Learning of Server Models in Model-Heterogeneous Federated learning

### v1

@@ Abstract
Federated learning (FL) is a distributed learning paradigm across multiple entities while preserving data privacy. However, with the continuous emergence of new data and increasing model diversity, traditional federated learning faces significant challenges, including inherent issues of data heterogeneity, model heterogeneity and catastrophic forgetting, along with new challenge of knowledge misalignment. In this study, we introduce FedDCL, a novel framework designed to enable data-free continual learning of the server model in a model-heterogeneous federated setting. We leverage pre-trained diffusion models to extract lightweight class-specific prototypes, which confer a threefold data-free advantage, enabling: (1) generation of synthetic data for the current task to augment training and counteract non-IID data distributions; (2) exemplar-free generative replay for retaining knowledge from previous tasks; and (3) data-free dynamic knowledge transfer from heterogeneous clients to the server. Experimental results on various datasets demonstrate the effectiveness of FedDCL, showcasing its potential to enhance the generalizability and practical applicability of federated learning in dynamic settings.
Figure 1: Illustration of the three main challenges in heterogeneous federated continual learning: (1) Catastrophic forgetting, where sequential task learning overwrites past knowledge and degrades earlier task performance; (2) Model heterogeneity, where diverse client architectures hinder direct parameter aggregation; and (3) Knowledge misalignment, where public dataset, due to static nature, fail to align with evolving task domains to transfer knowledge dynamically.
@@ I Introduction
In recent years, Federated Learning (FL) [37]has emerged as a machine learning paradigm that enables collaborative model training across multiple entities while preserving data privacy by keeping data localized. However, in real-world applications where data rapidly evolve, the classic FL framework faces significant challenges, such as catastrophic forgetting[36]. For better collecting, accumulating and leveraging knowledge from streaming data, Federated Continual Learning (FCL) has recently attracted growing attention [49, 58]. The FCL paradigm enables models to continuously adapt to new tasks or data distributions in the dynamic environment while retaining previously acquired knowledge. This capability has led to successful applications in multiple fields, such as action recognition[63], financial audit [40], person re-identification[68], and network traffic classification[52][29].
@@
In a typical FCL setting, the server generally assumes that all clients employ a completely homogeneous model architecture and communicates through parameter aggregation, as seen in methods like FedWeIT[60], FedKNOW[33], TARGET[64], and FedCIL[39]. However, with the widespread adoption of mobile and IoT devices, models have become increasingly diverse in architecture and capacity. For example, some large pretrained models deployed on cloud servers, such as BERT[9], ViT[11], and even modern large language models[44][1], often possess orders of magnitude more parameters than models on edge devices. This growing disparity in model scale and structure introduces significant model heterogeneity, which hinders effective collaboration towards obtaining a high-quality server model and ultimately limits the applicability and generalizability of FCL in real-world, heterogeneous environments.
@@
To bridge the barrier of architectural differences, Knowledge distillation[17] enables efficient knowledge exchange by communicating soft predictions. This approach has been adapted to the federated setting through Federated Distillation (FD), offering a compelling alternative to conventional parameter aggregation for knowledge exchange[21]. Methods like FedGems[5], Fed-ET[7], FedMKT[13], and FedBiOT[53] achieve heterogeneous model collaboration with an auxiliary dataset, bridging the model architecture gap. However, the above mentioned works are all limited to the single, static task. Although studies like FLwF-2T[46] and CFeD[34] combine distillation with dynamic settings in FCL, they primarily treat distillation as a regularization technique to mitigate forgetting, completely overlooking its potential for heterogeneous model collaboration. Therefore, the scenarios of leveraging heterogeneous models deployed across diverse clients to collectively and continuously train a larger server model in streaming knowledge remain an under-explored area.
@@
As shown in Figure. 1, heterogeneous federated continual learning faces three main challenges. First, the problem stems from catastrophic forgetting of server models in continual learning. As clients sequentially learn from non-stationary, streaming data over time, previously acquired knowledge tends to be overwritten or degraded, leading to performance collapse on earlier tasks. Then, model heterogeneity presents another major obstacle. Due to resource constraints, participating clients may deploy vastly different model architectures (e.g., varying depths, widths, or even model designs), which complicates knowledge transfer. Conventional federated parameter aggregation methods become ineffective under such architectural diversity. Finally, beyond the above inherent challenges posed by federated distillation and federated continual learning, their intersection introduces a novel challenge, which we term knowledge misalignment. The rapidly evolving data environment exacerbates the limitations of static public datasets commonly used in knowledge distillation. As illustrated in the toy example in Figure. 2, the effectiveness of distillation heavily depends on the domain alignment between the auxiliary dataset and the target task[43]. Manually select high-quality auxiliary datasets may expose sensitive domain information and compromising privacy. Moreover, in scenarios where domain knowledge evolves rapidly, the static nature of such datasets often struggles to accurately capture the dynamic changes in domain knowledge, thus hindering effective knowledge transfer [65]. Therefore, a practical problem arises: how can we enable heterogeneous device models to collaboratively and continuously learn a server model from streaming knowledge in a data-free manner ?
@@
Along this line, we propose FedDCL, a novel framework designed to facilitate knowledge collaboration among heterogeneous models to continuously train a server model under sequentially arriving tasks. We achieve data-free dynamic knowledge alignment in both knowledge transfer and knowledge retention. Specifically, FedDCL employs pre-trained diffusion models to generate compact, class-specific prototypes that capture the essential features of each class. This design offers three-fold benefits: (1) We can augment data for the current task through prototypes to strengthen training and mitigate non-IID challenges. (2) By replaying previously learned knowledge through synthesized samples and combining them with incoming real data during training, FedDCL effectively mitigates catastrophic forgetting and combats non-IID data shifts, all while avoiding the storage of any raw exemplars. (3) By adaptively generating synthetic data aligned with evolving task distributions, FedDCL enables seamless collaboration among heterogeneous clients, eliminating dependence on static public dataset. For the server model, we employ a multi-teacher distillation strategy, where knowledge from the latest task is transferred from clients and past models serve as teachers to mitigate catastrophic forgetting, effectively enabling the server model to continuously acquire knowledge. We conduct extensive experiments with various settings on two datasets, revealing the superiority of FedDCL over other baselines. Our code is available at: https://anonymous.4open.science/r/FCL-E7B9/. Our main contributions can be summarized as follows:
@@
  • We propose a novel federated continual learning framework that enables heterogeneous clients to collaboratively train a central server model under streaming tasks, overcoming the hybrid heterogeneity of models and data, catastrophic forgetting, and knowledge misalignment caused by continuously evolving task knowledge.
@@
  • We conduct extensive experiments with various settings across two datasets. Results reveal that FedDCL achieves a better performance, improving the server accuracy by 9.00%/6.48%/24.37%/18.96% on four settings.
@@ II Related Work
Federated learning is designed for privacy-sensitive scenarios, enabling collaborative training among decentralized clients without centralizing raw data. Classical federated learning algorithms, exemplified by FedAvg [37] and most of its variants[27][28], aggregate locally updated model parameters from participating devices to construct and refine a global server model. Despite the notable success of homogeneous federated learning, traditional strategies that assume uniform model architectures are becoming increasingly impractical as architectural heterogeneity grows more prevalent[59]. To tackle the bottleneck of training a global model with heterogeneous clients, existing model-heterogeneous federated learning approaches primarily follow two technical paradigms [12]: partial-training-based methods and knowledge-distillation-based methods. 1) Partial-training-based methods refer to approaches in which the server model is compressed to yield lightweight variants for localized training. Common approaches include quantization [38][35], which lowers numerical precision of weights, and pruning [10][22][69][51][6], which removes redundant parameters or structures. 2) Knowledge-distillation-based methods typically transfer knowledge using soft predictions or feature representations[17], without requiring architectural alignment.
@@
We focus on knowledge-distillation-based methods, which support higher degrees of model heterogeneity and are more common in practice. In this paradigm, clients train heterogeneous models locally and upload distilled knowledge. Methods like FedMD[26], DS-FL[20], FedDF[32], KT-pFL[66] and pFedHR[48] enable communication among heterogeneous clients through a proxy dataset. Additionally, as the model size on the server side continues to grow, some works like FedGems[5], FedKEMF[62], Fed-ET[7] and FedMKT[13] focus on utilizing clients to collaboratively train a larger server model. Most existing methods overly rely on a fixed public dataset that is closely aligned with the target domain to facilitate effective knowledge exchange. However, obtaining such a dataset is often unavailable in practice due to strict data privacy constraints. Although studies such as DaFKD [47] and PeFAD [55] do not rely on public datasets but instead generate synthetic data resembling the target distribution to enhance knowledge distillation, these data-free distillation methods are primarily designed to assist model-homogeneous federated learning. Besides, model-heterogeneous federated learning methods like FedGEN[70] FedGD[65] and FedZKT[67] train a global generator to produce synthetic data for replacing public data. However, training a generator from scratch is often time-consuming and the process may be unstable[3]. Last but not least, the above generative-based methods do not account for the dynamic matching process with continuous knowledge, rendering them ineffective in our context.
@@
Continual learning aims to train models sequentially on task flow while retaining previously acquired knowledge, a challenge commonly referred to as catastrophic forgetting[50]. Building upon this, Federated continual learning (FCL) has emerged as a promising paradigm that enables distributed clients to continuously collaborate while preserving privacy. Inspired by classic continual learning, FedWeIT[60] introduces the concept of federated continual learning. However, it focuses solely on the clients without considering the server model and necessitates the ID of the task for inference. Regularization-based methods, such as FLwF-2T[46] and CFeD[34] , mitigate catastrophic forgetting by penalizing parameter deviations through a knowledge distillation loss term. Data replay methods generally achieve stronger performance. Methods like FedKNOW[33] and FedViT[71] directly store real data from past tasks for replay, explicitly helping the model to recall previously learned knowledge. However, such direct storage of raw data may violate privacy constraints. Inspired by DGR[41], recent works achieve privacy-preserving data replay through synthetic data. Methods such as MFCL[2], TARGET[64], and FedCIL[39] employ adversarial training to learn generators that synthesize data for preserving previously acquired knowledge. Similarly, FedCBC[61] addresses spatial-temporal catastrophic forgetting via a VAE-based framework, yet it constructs separate binary classifiers for each class, limiting the flexibility. DDDR[31] leverages pre-trained diffusion models to perform class inversion for knowledge retention. It is worth noting that these methods are not designed to support collaboration among heterogeneous models. The recent work Mufti[14] explores continual collaboration among heterogeneous clients through the use of a public dataset. However, introducing such a dataset raises privacy leakage concerns. Moreover, it does not consider the server side model.
@@ III Problem Formulation
In this paper, we aim to continually train a server-side model through collaborative learning with heterogeneous client models over a sequence of tasks in a data-free manner. Specifically, we consider a federated learning system consisting of one central server and $K$ heterogeneous clients, indexed by $k=1,\dots,K$. Each client $k$ possesses a local model $\mathcal{M}_{k}$ , whose architecture may differ from others in depth, width, or structural design. Over time, all clients sequentially encounter a class-incremental task flow $\{t\}_{t=1}^{T}$, where each task $t$ introduces a new set of classes that are strictly disjoint from those in all previous and future tasks. Within each task $t$, client $k$ holds a private local dataset $\mathcal{D}_{k}^{(t)}=\{(x_{i}^{(t)},y_{i}^{(t)})\}_{i=1}^{N_{k}^{(t)}}$, where $N_{k}^{(t)}$ denotes the number of samples available to client $k$ at task $t$. The data distributions $\{\mathcal{D}_{k}^{(t)}\}_{k=1}^{K}$ are typically non-IID across clients.
The objective is to effectively train the server model $\mathcal{M}_{s}$ using heterogeneous client models, enabling it to accumulate knowledge across all tasks while avoiding catastrophic forgetting. This could be formalized as:
@@ IV METHODOLOGY
Given the continuous changes in the target domain across tasks, it is crucial to dynamically align with emerging knowledge and save replayable representations in a storage-efficient manner. Inspired by class inversion based on pre-trained diffusion models[15] and its recent success in the federated setting[31][56][57], we further investigate its potential in the model-heterogeneous federated continual learning, a largely unexplored yet critical setting. Building upon this technique, we implement an effective knowledge extraction mechanism to efficiently align and preserve category-level information across the continual task flow.
Diffusion Process. Diffusion models are a class of generative models that learn to synthesize data by reversing a gradual noising process[18][42]. We denote $p_{\text{data}}$ as the raw data distribution and $p_{\text{latent}}=\mathcal{N}(\mathbf{0},\mathbf{I})$ as the target latent distribution, where $\mathcal{N}$ is the Gaussian distribution. The forward diffusion process follows a fixed Markov chain that progressively adds noise to original distribution $p_{\text{data}}$ until it converges to $p_{\text{latent}}$. During the reverse denoising process, the model is trained to iteratively remove the noise to construct the desired data. Formally, let $x_{0}\sim p_{\text{data}}$. The whole forward process can be formulated as:

### v2

# Data-Free Continual Learning of Server Models in Model-Heterogeneous Cloud-Device Collaboration

@@ Abstract
The rise of cloud-device collaborative computing has enabled intelligent services to be delivered across distributed edge devices while leveraging centralized cloud resources. In this paradigm, federated learning (FL) has become a key enabler for privacy-preserving model training without transferring raw data from edge devices to the cloud. However, with the continuous emergence of new data and increasing model diversity, traditional federated learning faces significant challenges, including inherent issues of data heterogeneity, model heterogeneity and catastrophic forgetting, along with new challenge of knowledge misalignment. In this study, we introduce FedDCL, a novel framework designed to enable data-free continual learning of the server model in a model-heterogeneous federated setting. We leverage pre-trained diffusion models to extract lightweight class-specific prototypes, which confer a threefold data-free advantage, enabling: (1) generation of synthetic data for the current task to augment training and counteract non-IID data distributions; (2) exemplar-free generative replay for retaining knowledge from previous tasks; and (3) data-free dynamic knowledge transfer from heterogeneous devices to the cloud server. Experimental results on various datasets demonstrate the effectiveness of FedDCL, showcasing its potential to enhance the generalizability and practical applicability of federated cloud-device collaboration in dynamic settings.
Figure 1: Illustration of the three main challenges in heterogeneous federated continual learning: (1) Catastrophic forgetting, where sequential task learning overwrites past knowledge and degrades earlier task performance; (2) Model heterogeneity, where diverse device architectures hinder direct parameter aggregation; and (3) Knowledge misalignment, where public dataset, due to static nature, fail to align with evolving task domains to transfer knowledge dynamically.
@@ I Introduction
In recent years, the proliferation of edge devices, such as smartphones, IoT sensors, and surveillance cameras, has driven a paradigm shift toward cloud-device collaborative computing [Cloud_edge_intro]. Within this ecosystem, Federated Learning (FL) [Fedavg] has emerged as a representative machine learning paradigm that enables collaborative model training across multiple entities while preserving data privacy by keeping data localized. This design is strongly motivated by stringent data protection regulations, such as the GDPR [gdpr] and CCPA [ccpa], which impose strict limitations on the centralized collection and processing of personal data.
@@
However, in real-world applications where data rapidly evolve, the classic FL framework faces significant challenges, such as catastrophic forgetting[catastrophic_forgetting]. For better collecting, accumulating and leveraging knowledge from streaming data, federated continual learning (FCL) has recently attracted growing attention [FCL_survey1][FCL_survey2]. The FCL paradigm enables models to continuously adapt to new tasks or data distributions in the dynamic environment while retaining previously acquired knowledge. This capability has led to successful applications in multiple fields, such as action recognition[application_ar], financial audit [application_fin], person re-identification[application_id], and network traffic classification[trafic][trafic2]. In a typical FCL setting, the server generally assumes that all devices employ a completely homogeneous model architecture and communicates through parameter aggregation, as seen in methods like FedWeIT[Fedwelt], FedKNOW[fedknow], TARGET[target], and FedCIL[fedcil]. However, with the widespread adoption of mobile and IoT devices, models have become increasingly diverse in architecture and capacity. For example, some large pretrained models deployed on cloud servers, such as BERT[bert], ViT[vit], and even modern large language models[touvron2023llama][achiam2023gpt], often possess orders of magnitude more parameters than models on edge devices. This growing disparity in model scale and structure introduces significant model heterogeneity, which hinders effective collaboration towards obtaining a high-quality server model and ultimately limits the applicability and generalizability of FCL in real-world, heterogeneous environments.
@@
To bridge the barrier of architectural differences, knowledge distillation[hinton2015distilling] enables efficient knowledge exchange by communicating soft predictions. This approach has been adapted to the federated setting through federated distillation (FD), offering a compelling alternative to conventional parameter aggregation for knowledge exchange[FD]. Methods like FedGems[fedgems], Fed-ET[Fedet], FedMKT[fan2024fedmkt], and FedBiOT[wu2024fedbiot] achieve heterogeneous model collaboration with an auxiliary dataset, bridging the model architecture gap. However, the above mentioned works are all limited to the single, static task. Although studies like FLwF-2T[fedLWF2T] and CFeD[ma2022continual] combine distillation with dynamic settings in FCL, they primarily treat distillation as a regularization technique to mitigate forgetting, completely overlooking its potential for heterogeneous model collaboration. Therefore, the scenarios of leveraging heterogeneous models deployed across diverse devices to collectively and continuously train a larger server model in streaming knowledge remain an under-explored area.
@@
As shown in Figure. 1, heterogeneous federated continual learning faces three main challenges. First, the problem stems from catastrophic forgetting of server models in continual learning. As clients sequentially learn from non-stationary, streaming data over time, previously acquired knowledge tends to be overwritten or degraded, leading to performance collapse on earlier tasks. Then, model heterogeneity presents another major obstacle. Due to resource constraints, participating devices may deploy vastly different model architectures (e.g., varying depths, widths, or even model designs), which complicates knowledge transfer. Conventional federated parameter aggregation methods become ineffective under such architectural diversity. Finally, beyond the above inherent challenges posed by federated distillation and federated continual learning, their intersection introduces a novel challenge, which we term knowledge misalignment. The rapidly evolving data environment exacerbates the limitations of static public datasets commonly used in knowledge distillation. As illustrated in the toy example in Figure. 2, the effectiveness of distillation heavily depends on the domain alignment between the auxiliary dataset and the target task[su2022domain]. Manually select high-quality auxiliary datasets may expose sensitive domain information and compromising privacy. Moreover, in scenarios where domain knowledge evolves rapidly, the static nature of such datasets often struggles to accurately capture the dynamic changes in domain knowledge, thus hindering effective knowledge transfer [zhang2023towards]. Therefore, a practical problem arises: how can we enable heterogeneous device models to collaboratively and continuously learn a server model from streaming knowledge in a data-free manner?
@@
Along this line, we propose FedDCL, a novel framework designed to facilitate knowledge collaboration among heterogeneous models to continuously train a server model under sequentially arriving tasks. We achieve data-free dynamic knowledge alignment in both knowledge transfer and knowledge retention. Specifically, FedDCL employs pre-trained diffusion models to generate compact, class-specific prototypes that capture the essential features of each class. This design offers three-fold benefits: (1) We can augment data for the current task through prototypes to strengthen training and mitigate non-IID challenges. (2) By replaying previously learned knowledge through synthesized samples and combining them with incoming real data during training, FedDCL effectively mitigates catastrophic forgetting and combats non-IID data shifts, all while avoiding the storage of any raw exemplars. (3) By adaptively generating synthetic data aligned with evolving task distributions, FedDCL enables seamless collaboration among heterogeneous devices, eliminating dependence on static public dataset. For the server model, we employ a multi-teacher distillation strategy, where knowledge from the latest task is transferred from edge clients and past models serve as teachers to mitigate catastrophic forgetting, effectively enabling the server model to continuously acquire knowledge. We conduct extensive experiments with various settings on two datasets, revealing the superiority of FedDCL over other baselines. Our code is available at: https://anonymous.4open.science/r/FCL-E7B9. Our main contributions can be summarized as follows:
@@
  • We propose a novel federated continual learning framework that enables heterogeneous devices to collaboratively train a central server model under streaming tasks, overcoming the hybrid heterogeneity of models and data, catastrophic forgetting, and knowledge misalignment caused by continuously evolving task knowledge.
@@
  • We conduct extensive experiments with various settings across two datasets. Results reveal that FedDCL achieves a better performance, improving the server accuracy by 9.00%/6.48%/23.67%/16.40% on four settings.
@@ II Related Work
Federated learning is designed for privacy-sensitive scenarios, enabling collaborative training among decentralized clients without centralizing raw data. Classical federated learning algorithms, exemplified by FedAvg [Fedavg] and most of its variants[FedMoon][FedProx], aggregate locally updated model parameters from participating devices to construct and refine a global server model. Despite the notable success of homogeneous federated learning, traditional strategies that assume uniform model architectures are becoming increasingly impractical as architectural heterogeneity grows more prevalent[heterogeneous_survey1]. To tackle the bottleneck of training a global model with heterogeneous clients, existing model-heterogeneous federated learning approaches primarily follow two technical paradigms [heterogeneous_survey2]: partial-training-based methods and knowledge-distillation-based methods. 1) Partial-training-based methods refer to approaches in which the server model is compressed to yield lightweight variants for localized training. Common approaches include quantization [ozkara2021quped][markov2023quantized], which lowers numerical precision of weights, and pruning [diao2020heterofl][jiang2022model][zhu2022resilient][wang2023theoretical][cho2024heterogeneous], which removes redundant parameters or structures. 2) Knowledge-distillation-based methods typically transfer knowledge using soft predictions or feature representations[hinton2015distilling], without requiring architectural alignment.
@@
We focus on knowledge-distillation-based methods, which support higher degrees of model heterogeneity and are more common in practice. In this paradigm, clients train heterogeneous models locally and upload distilled knowledge. Methods like FedMD[li2019fedmd], DS-FL[itahara2021distillation], FedDF[lin2020ensemble], KT-pFL[KT-pFL] and pFedHR[pFedHR] enable communication among heterogeneous clients through a proxy dataset. Additionally, as the model size on the server side continues to grow, some works like FedGems[fedgems], FedKEMF[FedKEMF], Fed-ET[Fedet] and FedMKT[fan2024fedmkt] focus on utilizing clients to collaboratively train a larger server model. Most existing methods overly rely on a fixed public dataset that is closely aligned with the target domain to facilitate effective knowledge exchange. However, obtaining such a dataset is often unavailable in practice due to strict data privacy constraints. Although studies such as DaFKD [wang2023dafkd] and PeFAD [xu2024pefad] do not rely on public datasets but instead generate synthetic data resembling the target distribution to enhance knowledge distillation, these data-free distillation methods are primarily designed to assist model-homogeneous federated learning. Besides, model-heterogeneous federated learning methods like FedGEN[zhu2021data] FedGD[zhang2023towards] and FedZKT[fedzkt] train a global generator to produce synthetic data for replacing public data. However, training a generator from scratch is often time-consuming and the process may be unstable[gan_collapse]. Last but not least, the above generative-based methods do not account for the dynamic matching process with continuous knowledge, rendering them ineffective in our context.
@@
Continual learning aims to train models sequentially on task flow while retaining previously acquired knowledge, a challenge commonly referred to as catastrophic forgetting[cl_survey]. Building upon this, Federated continual learning (FCL) has emerged as a promising paradigm that enables distributed clients to continuously collaborate while preserving privacy. Inspired by classic continual learning, FedWeIT[Fedwelt] introduces the concept of federated continual learning. However, it focuses solely on the clients without considering the server model and necessitates the ID of the task for inference. Regularization-based methods, such as FLwF-2T[fedLWF2T] and CFeD[ma2022continual] , mitigate catastrophic forgetting by penalizing parameter deviations through a knowledge distillation loss term. Data replay methods generally achieve stronger performance. Methods like FedKNOW[fedknow] and FedViT[fedvit] directly store real data from past tasks for replay, explicitly helping the model to recall previously learned knowledge. However, such direct storage of raw data may violate privacy constraints. Inspired by DGR[shin2017continual], recent works achieve privacy-preserving data replay through synthetic data. Methods such as MFCL[babakniya2024data], TARGET[target], and FedCIL[fedcil] employ adversarial training to learn generators that synthesize data for preserving previously acquired knowledge. Similarly, FedCBC[yu2024overcoming] addresses spatial-temporal catastrophic forgetting via a VAE-based framework, yet it constructs separate binary classifiers for each class, limiting the flexibility. DDDR[dddr] leverages pre-trained diffusion models to perform class inversion for knowledge retention. It is worth noting that these methods are not designed to support collaboration among heterogeneous models. The recent work Mufti[mufti] explores continual collaboration among heterogeneous clients through the use of a public dataset. However, introducing such a dataset raises privacy leakage concerns. Moreover, it does not consider the server side model.
@@ III Problem Formulation
In this paper, we aim to continually train a server-side model through collaborative learning with heterogeneous device models over a sequence of tasks in a data-free manner. Specifically, we consider a federated learning system consisting of one central server and $K$ heterogeneous clients, indexed by $k=1,\dots,K$. Each client $k$ possesses a local model $\mathcal{M}_{k}$ , whose architecture may differ from others in depth, width, or structural design. Over time, all clients sequentially encounter a class-incremental task flow $\{t\}_{t=1}^{T}$, where each task $t$ introduces a new set of classes that are strictly disjoint from those in all previous and future tasks. Within each task $t$, client $k$ holds a private local dataset $\mathcal{D}_{k}^{(t)}=\{(x_{i}^{(t)},y_{i}^{(t)})\}_{i=1}^{N_{k}^{(t)}}$, where $N_{k}^{(t)}$ denotes the number of samples available to client $k$ at task $t$. The data distributions $\{\mathcal{D}_{k}^{(t)}\}_{k=1}^{K}$ are typically non-IID across clients.
The objective is to effectively train the server model $\mathcal{M}_{s}$ using heterogeneous models, enabling it to accumulate knowledge across all tasks while avoiding catastrophic forgetting. This could be formalized as:
@@ IV METHODOLOGY
Given the continuous changes in the target domain across tasks, it is crucial to dynamically align with emerging knowledge and save replayable representations in a storage-efficient manner. Inspired by class inversion based on pre-trained diffusion models[textual_inversion] and its recent success in the federated setting[dddr][yang2024exploring][yang2024feddeo], we further investigate its potential in the model-heterogeneous federated continual learning, a largely unexplored yet critical setting. Building upon this technique, we implement an effective knowledge extraction mechanism to efficiently align and preserve category-level information across the continual task flow.
Diffusion Process. Diffusion models are a class of generative models that learn to synthesize data by reversing a gradual noising process[ddpm][ddim]. We denote $p_{\text{data}}$ as the raw data distribution and $p_{\text{latent}}=\mathcal{N}(\mathbf{0},\mathbf{I})$ as the target latent distribution, where $\mathcal{N}$ is the Gaussian distribution. The forward diffusion process follows a fixed Markov chain that progressively adds noise to original distribution $p_{\text{data}}$ until it converges to $p_{\text{latent}}$. During the reverse denoising process, the model is trained to iteratively remove the noise to construct the desired data. Formally, let $x_{0}\sim p_{\text{data}}$. The whole forward process can be formulated as: